# Search for baryon junctions in e+A collisions at the Electron Ion Collider

### v1

@@ I INTRODUCTION
In the context of ultra-relativistic nucleus-nucleus collisions, a substantial puzzle has been the observed baryon asymmetry at mid-rapidity, a phenomenon inconsistent with the expected transport behavior of valence quarks Gyulassy et al. (1997); Appelshauser et al. (1999); Arsene et al. (2009); Abelev et al. (2009); Adamczyk et al. (2017); Aamodt et al. (2010). The expected behavior follows from the fact that valence quarks, carrying a large fraction of the baryon momentum and therefore having shorter interaction times, should end up in the fragmentation regions far from mid-rapidity Kharzeev (1996). Nevertheless, the baryon junction model offers a compelling explanation for this discrepancy. Baryon junctions, composed of low-momentum gluons, can interact with the soft parton field of the projectile and become stopped at mid-rapidity. This mechanism leads to the production of baryons with low transverse momentum and different quark content than the colliding baryons and accompanying pions, leading to high multiplicity events. In addition, the exponential drop in cross-section with rapidity loss, characteristic of junction-stopping processes, provides a distinctive signature that supports the junction model over the valence quark model in explaining baryon transport in high-energy collisions. For example, in a semi-inclusive photon-induced process off a proton such as $\gamma^{*}+p\rightarrow B+X$, the rapidity distribution of the produced net-baryon is expected to follow
@@
In the 2030s, the Electron-Ion Collider (EIC) is set to advance nuclear physics research by supplying nuclear deep inelastic scattering with collider kinematics Abdul Khalek et al. (2022). In the context of e+A collisions at the EIC, the reaction mechanism can be divided into three stages Chang et al. (2022): (i) hard scattering processes Piller et al. (1995), (ii) IntraNuclear Cascade (INC) processes Cugnon (1982), and (iii) the breakup of excited nuclear remnants Weisskopf (1937). Isolating kinematic regions dominated by each process will be essential for refining nuclear excitation models and gaining deeper insights into the target nucleus Mathews et al. (1982); Magdy et al. (2024). Therefore, our work addresses the previously raised question in the EIC context and within the framework of the Benchmark eA Generator for Leptoproduction (BeAGLE) model.
@@ II MODEL AND METHODOLOGY
Figure 2: Schematic representation of the key quantities describing the collision geometry, including $b$, the impact parameter; $R$, the spatial displacement from the interaction point to the center of the nucleus; and $d$, the projected virtual photon travel distance from the interaction point to the edge of the nuclear medium.
@@ III RESULTS
Figure 3: Net-baryon rapidity distributions from the PYTHIA model for (i) all processes included and process 99. The solid line represents the fit using Eq. 1.
The net-baryon distributions for e+p collisions at $10\times 40$ GeV from the PYTHIA model are illustrated in Fig. 3. Our findings suggest that the dominant contributor to the net-baryon distribution is the PYTHIA process of LO DIS (Process 99). The results exhibit a linear relationship between the net-baryon and $y-Y_{beam}$ with a slope of $\alpha_{J}=1.5$. It is worth noting that the net-baryon distribution from either the quark model scenario or the string junction scenario is anticipated to differ only in the slope $\alpha_{J}$. Therefore, the extracted slope of $\alpha_{J}=1.5$ exceeds the theoretical expectation Kharzeev (1996) for the baryon junction, where $\alpha_{J}=-0.5,0,0.5$, is inline with the absence of baryon junction physics in the PYTHIA model utilized in this study.
Figure 4: Net-baryon rapidity distributions from the BeAGLE model for PYTHIA processes panel (a) and BeAGLE processes panel (b).
In contrast to the straightforward depiction presented in the e+p collisions 3, the measurements of net-baryon distribution in e+A collisions are anticipated to be more intricate. This anticipation stems from our understanding that the reaction mechanism in e+A collisions at the EIC encompasses various processes, including (i) hard scattering Piller et al. (1995), (ii) IntraNuclear Cascade Cugnon (1982), and (iii) breakup of excited nuclear remnants Weisskopf (1937). Therefore, it is crucial to expand our examination of the net-baryon distribution to include e+A collisions. Figure 4 displays the net-baryon distributions for e+Au collisions at $10\times 40$ GeV from the BeAGLE model for multiple PYTHIA processes (a) and several BeAGLE processes (b). The net-baryon distribution results in panel (a) demonstrate varying dependencies on $y-Y_{beam}$. This variability arises from the different characteristics of particle production dependence on $y-Y_{beam}$ in the BeAGLE model. The extracted slope $\alpha_{J}$ at different $y-Y_{beam}$ values differs and exceeds the theoretical expectation for the baryon junction Kharzeev (1996). These findings are in line with the absence of baryon junction physics in the BeAGLE model. Furthermore, measuring the net-baryon distribution for e+(p)Au collisions could serve as one of the initial EIC measurements to verify the validation of the quark model scenario.
@@
The EMC Collaboration previously measured proton transport in $\mu+p$ vs. $\mu+d$ Arneodo et al. (1987). Figure 7 presents a modified plot of these measurements, with data extracted from Ref Arneodo et al. (1987). The $W$ dependence of the net-proton for $\mu$+p and $\mu$+d at 280-GeV muon on fixed targets is presented in Fig. 7 panel (a). The EMC measurements indicated a weak dependence on the target flavor. In contrast, the LUND and BeAGLE models’ calculations suggested about 20% higher proton yield for the proton target compared to the deuteron target. Additionally, panel (b) of Fig. 7 shows the ratio of net-proton yields between $\mu$+p and $\mu$+d from both data and simulations. The results in panel (b) indicated that while the data is target flavor independent (within the uncertainties), the simulations from the LUND and BeAGLE models revealed approximately 20% target flavor dependence. Consequently, the EMC Collaboration’s measurements suggest to be inconsistent with the naive expectation that proton yield would correspond to the target flavor, meaning more protons from the proton target than the deuteron target if the baryon number is directly related to valence quarks. The EMC Collaboration stated that:“ Proton as well as antiproton production from both a hydrogen and a deuterium target are found to show the same behaviour in the $x_{F}$ distributions. This target independence is also true for the $W$ dependence of the average multiplicities. The multiplicities of antiprotons and forward produced protons exhibit the same significant rise with $W$. The Lund model is generally in agreement with the data. However it predicts a somewhat higher yield of protons from the hydrogen than from the deuterium target, an effect which is less pronounced in the data.” Arneodo et al. (1987). Additionally, our BeAGLE model results align with the Lund model’s conclusions. Therefore, more detailed measurements of the net-proton yields, with high statistics and different isobar targets at the future EIC, would help determine whether proton yields at given kinematics are correlated with the quark flavor of the target.
@@ IV CONCLUSION
In summary, we investigate the potential answer to the question, “What traces the baryon number?” Kharzeev (1996) at the future EIC. We discussed (i) the Regge intercept of the net-baryon distributions in e+(p)Au collisions, (ii) the isobaric ratio constructed from the e+Ru and e+Zr collisions, and (iii) the EMC net-proton ratios at the future EIC. Our results for the extracted Regge intercept in e+(p)Au collisions indicated a value consistent with the valence-quarks scenario. The $R(Isobar)$ showed values less than unity and independent of $x_{B_{j}}$, $Q^{2}$, and momentum. Additionally, the EMC net-proton ratios challenge the expectation that proton yield would follow the target flavor in the model simulation. These simulations are consistent with the valence quarks carrying the baryon number, as included in the model we used, but appear to be inconsistent with experimental results Lewis et al. (2024); Arneodo et al. (1987) from the STAR and EMC collaborations. We conclude that conducting detailed measurements of the net-baryon distributions in e+(p)Au collisions, the isobaric ratio, and the EMC net-proton ratios at the future EIC can help determine what carries the baryon number.

### v2

@@ I INTRODUCTION
In the context of ultra-relativistic nucleus-nucleus collisions, a substantial puzzle has been the observed baryon asymmetry at mid-rapidity, a phenomenon inconsistent with the expected transport behavior of valence quarks Gyulassy et al. (1997); Appelshauser et al. (1999); Arsene et al. (2009); Abelev et al. (2009); Adamczyk et al. (2017); Aamodt et al. (2010). The expected behavior follows from the fact that valence quarks, carrying a large fraction of the baryon momentum and therefore having shorter interaction times, should end up in the fragmentation regions far from mid-rapidity Kharzeev (1996); Cebra et al. (2022). Nevertheless, the baryon junction model offers a compelling explanation for this discrepancy. Baryon junctions, composed of low-momentum gluons, can interact with the soft parton field of the projectile and become stopped at mid-rapidity. This mechanism leads to the production of baryons with low transverse momentum and different quark content than the colliding baryons and accompanying pions, leading to high multiplicity events. In addition, the exponential drop in cross-section with rapidity loss, characteristic of junction-stopping processes, provides a distinctive signature that supports the junction model over the valence quark model in explaining baryon transport in high-energy collisions. For example, in a semi-inclusive photon-induced process off a proton such as $\gamma^{*}+p\rightarrow B+X$, the rapidity distribution of the produced net-baryon is expected to follow
@@
In the 2030s, the Electron-Ion Collider (EIC) is set to advance nuclear physics research by supplying nuclear deep inelastic scattering with collider kinematics Abdul Khalek et al. (2022). In the context of e+A collisions at the EIC, the reaction mechanism can be categorized into three distinct stages Chang et al. (2022): (i) hard scattering processes Piller et al. (1995), (ii) intra-nuclear cascade (INC) processes Cugnon (1982), and (iii) the breakup of excited nuclear remnants Weisskopf (1937). Identifying kinematic regions dominated by each stage is crucial for refining nuclear excitation models and deepening our understanding of the target nucleus Mathews et al. (1982); Magdy et al. (2024). Ref. Cebra et al. (2022) identifies backward ($u$-channel) production at the EIC as a promising approach for exploring baryon number transport mechanisms, potentially connected to baryon junction dynamics. Our work addresses the three suggested measurements within the EIC framework, utilizing the Benchmark eA Generator for Leptoproduction (BeAGLE) model.
@@ II MODEL AND METHODOLOGY
Figure 2: The top panel gives a schematic representation of the quantities describing the collision geometry, including $b$, the impact parameter; $R$, the spatial displacement from the interaction point to the center of the nucleus; and $d$, the projected virtual photon travel distance from the interaction point to the edge of the nuclear medium. The lower panel displays the $dN/d\eta$ distribution of the $e$+$Au$ collisions at 10$\times$40 (GeV) from the BeAGLE model. The shaded areas represent the three $\eta$ acceptance of the three mean systems expected by the ePIC experiment at the EIC Abdul Khalek et al. (2022); ePI as (i) $-3.5<\eta<3.5$, (ii) $5.94.5<\eta<$ 5.9 and (iii) $\eta>6.0$.
The analysis focuses on $e+p(A)$ collisions at 10$\times$40 GeV, with approximately 0.3 billion events generated for each scenario considered. With the anticipated luminosity of the EIC, around 400 billion events are expected, allowing our analyses to be conducted with sufficient statistical power at the EIC.
@@ III RESULTS
Figure 3: Net-baryon rapidity distributions of e+p collisions at 10$\times$40 GeV from the PYTHIA model for all processes included and process 99 panel (a), and different $p_{T}$ cuts panel (b). The solid line represents the fit using Eq. 1. The dashed line gives the expectation from a junction scenario Kharzeev (1996).
The net-baryon distributions for e+p collisions at $10\times 40$ GeV from the PYTHIA model are illustrated in Fig. 3. Our findings suggest that the dominant contributor to the net-baryon distribution is the PYTHIA process of LO DIS (Process 99). The results exhibit a linear relationship between the net-baryon and $y-Y_{beam}$ with a slope of $\alpha_{J}=1.5$. It is worth noting that the net-baryon distribution from either the quark model scenario or the string junction scenario is anticipated to differ only in the slope $\alpha_{J}$. Therefore, the extracted slope of $\alpha_{J}=1.5$ exceeds the theoretical expectation Kharzeev (1996) for the baryon junction, where $\alpha_{J}=0.5$ (dashed line), and is inline with the absence of baryon junction physics in the PYTHIA model utilized in this study. Additionally, we present the net-baryon distributions for two $p_{T}$ cuts. Our results show minimal differences, if any, between the no-$p_{T}$ cut and $p_{T}<0.1\,\text{GeV}/c$. It is important to emphasize that hadron yields can be measured with an extrapolation of $p_{T}$ to zero rather than being limited to the detector’s $p_{T}$ acceptance when compared to thermal models STA (2024); Adamczyk et al. (2017). While uncertainties are inherent in such extrapolations, we should present a physical quantity independent of specific detector acceptance and configuration to compare to theory and other measurements.
Figure 4: Net-baryon rapidity distributions of e+Au collisions at 10$\times$40 GeV from the BeAGLE model for PYTHIA processes panel (a) and BeAGLE processes panel (b).
In contrast to the straightforward depiction presented in the e+p collisions 3, the measurements of net-baryon distribution in e+A collisions are anticipated to be more intricate. This anticipation stems from our understanding that the reaction mechanism in e+A collisions at the EIC encompasses various processes, including (i) hard scattering Piller et al. (1995), (ii) IntraNuclear Cascade Cugnon (1982), and (iii) breakup of excited nuclear remnants Weisskopf (1937) (see lower pane of Fig. 2). Therefore, expanding our examination of the net-baryon distribution to e+A collisions is crucial to understand its role in our baryon stooping study. Figure 4 displays the net-baryon distributions for e+Au collisions at $10\times 40$ GeV from the BeAGLE model for multiple PYTHIA processes (a) and several BeAGLE processes (b). The net-baryon distribution results in panel (a) demonstrate varying dependencies on $y-Y_{beam}$. This variability arises from the different characteristics of particle production dependence on $y-Y_{beam}$ in the BeAGLE model given in panel (b). Consequently, interpreting the results of the net-baryon distribution for e+Au collisions is not straightforward. Although the slope, $\alpha_{J}$, at various values of $y-Y_{beam}$ differs and exceeds theoretical expectations for the baryon junction Kharzeev (1996), further theoretical insight is needed to fully understand the baryon junction’s expected behavior in e+A collisions.
@@
The EMC Collaboration previously measured proton transport in $\mu+p$ vs. $\mu+d$ Arneodo et al. (1987). Figure 7 presents a modified plot of these measurements, with data extracted from Ref Arneodo et al. (1987). The $W$ dependence of the net-proton for $\mu$+p and $\mu$+d at 280-GeV muon on fixed targets is presented in Fig. 7 panel (a). The EMC measurements indicated a weak dependence on the target flavor. In contrast, the LUND and BeAGLE models’ calculations suggested about 20% higher proton yield for the proton target compared to the deuteron target. Additionally, panel (b) of Fig. 7 shows the ratio of net-proton yields between $\mu$+p and $\mu$+d from both data and simulations. The results in panel (b) indicated that while the data is target flavor independent (within the uncertainties), the simulations from the LUND and BeAGLE models revealed approximately 20% target flavor dependence.
Consequently, the EMC Collaboration’s measurements suggest to be inconsistent with the naive expectation that proton yield would correspond to the target flavor, meaning more protons from the proton target than the deuteron target if the baryon number is directly related to valence quarks. The EMC Collaboration stated that:“ Proton as well as antiproton production from both a hydrogen and a deuterium target are found to show the same behaviour in the $x_{F}$ distributions. This target independence is also true for the $W$ dependence of the average multiplicities. The multiplicities of antiprotons and forward produced protons exhibit the same significant rise with $W$. The Lund model is generally in agreement with the data. However it predicts a somewhat higher yield of protons from the hydrogen than from the deuterium target, an effect which is less pronounced in the data.” Arneodo et al. (1987). Additionally, our BeAGLE model results align with the Lund model’s conclusions. Therefore, more detailed measurements of the net-proton yields, with high statistics and different isobar targets at the future EIC, would help determine whether proton yields at given kinematics are correlated with the quark flavor of the target.
@@ IV CONCLUSION
In summary, we investigate the potential answer to the question, “What traces the baryon number?” Kharzeev (1996) at the future EIC. We discussed (i) the Regge intercept of the net-baryon distributions in e+(p)Au collisions, (ii) the isobaric ratio constructed from the e+Ru and e+Zr collisions, and (iii) the EMC net-proton ratios at the future EIC. Our results for the extracted Regge intercept in e+(p)Au collisions indicated a value consistent with the valence-quarks scenario. The $R(Isobar)$ showed values less than unity and independent of $x_{B_{j}}$, $Q^{2}$, and momentum. Additionally, the EMC net-proton ratios challenge the expectation that proton yield would follow the target flavor in the model simulation. These simulations are consistent with the valence quarks carrying the baryon number, as included in the model we used, but appear to be inconsistent with experimental results Lewis et al. (2024); Arneodo et al. (1987) from the STAR and EMC collaborations. We conclude that conducting detailed measurements of the net-baryon distributions in e+(p)Au collisions, the isobaric ratio and the EMC net-proton ratios at the future EIC can help determine what carries the baryon number.